OpenStudy (xenesis98):

Robin randomly selects a number between 1 and 20. What is the probability that the number selected is the square of a natural number? a.1/2 b.-3/20 c.-3/10 d.1/5 e.1/3

OpenStudy (bobo-i-bo):

P(Square number selected)=Number of square numbers / Total number of numbers

OpenStudy (xenesis98):

so?

OpenStudy (bobo-i-bo):

It's against Open Study to tell you answers I'm afraid. But i can try and help you :) Did you understand what i just said?

OpenStudy (xenesis98):

nope

OpenStudy (debbieg):

How man numbers between 1 and 20 are squares of natural numbers?

OpenStudy (debbieg):

*many

OpenStudy (xenesis98):

what are natural numbers

OpenStudy (r-o-s-e):

natraul numbers are whole numbers

OpenStudy (r-o-s-e):

do you get it now

OpenStudy (xenesis98):

so how do you find out

OpenStudy (debbieg):

You list the numbers between 1 and 20 that are perfect squares. How many are there?

OpenStudy (xenesis98):

you mean like 4,9,16?

OpenStudy (r-o-s-e):

something like that but, i would say yes

OpenStudy (xenesis98):

so is it 3?

OpenStudy (xenesis98):

3 squares?

OpenStudy (r-o-s-e):

i would say so

OpenStudy (debbieg):

No, you are forgetting one. Literally. ;)

OpenStudy (debbieg):

Also, 3/20 is not an answer choice. So that should be a red flag.

OpenStudy (xenesis98):

oh true so itll 1/5

OpenStudy (r-o-s-e):

wait a sec

OpenStudy (r-o-s-e):

it most likely is 1/5 now i got to go

OpenStudy (xenesis98):

ok

OpenStudy (debbieg):

Yes, 4 squares out of 20 numbers = 4/20 = 1/5